kekeman:

Which of the following expressions could not be used to determine the exact value of cos 120°? cos2(60°) − sin2(60°) 2cos2(60°) − 1 2sin(60°)cos(60°) 1 − 2sin2(60°)

jhonyy9:

try it in this way 120 = 60 + 60 cos 120° = cos (60+60) = ? use formula cos (a+b)= ?

kekeman:

@jhonyy9 wrote:
try it in this way 120 = 60 + 60 cos 120° = cos (60+60) = ? use formula cos (a+b)= ?
I think the answer is 2sin(60°)cos(60°)